Amazon has a "Lightning Deal" where it slashes the price of one item. At 3:15 p.m. today, they announced that the item was a new tablet with a sales price of $150$150 which is less than half the original price.Amazon has a Lightning Deal where it slashes the price of one item. At 3:15 p.m. today, they announced that the item was a
Answer #1

When the price is $150, only first four students will purchase the item as willingness to pay of fifth student < 150.
So, Total consumer surplus = (500 - 150) + (400 - 150) + (300 - 150) + (200 - 150) = 350 + 250 + 150 + 50 = $800

When the price is $350, only first two students will purchase the item as willingness to pay for others < 350.
So, Total consumer surplus = (500 - 350) + (400 - 350) = 150 + 50 = $200

So, change in total consumer surplus = 200 - 800 = -$600